OpenStudy (anonymous):

Convert (-1,square root 3) into polar coordinates.

OpenStudy (anonymous):

\[(-1,\sqrt{3})\]

OpenStudy (anonymous):

r=(1+3)^(1/2), theta= 2Pi/3